OpenStudy (anonymous):

for each x belonging to [0,1], let f(x) = x if x is rational, and let f(x) = 1-x if x is irrational. Prove that f(f(x)) = x for all x belonging [0,1].

OpenStudy (amistre64):

so its a piecewise function eh

OpenStudy (amistre64):

f(x) = x, or 1-x let x = a f(f(a)) = ((a)), or (1-(a)), or 1-(1-a)

OpenStudy (amistre64):

maybe defining the parts of f in their own right might make this format easier to read

OpenStudy (amistre64):

r(x) = x; i(x) = 1-x r(r) = x r(i) = 1-x i(i) = 1-(1-x) = x i(r) = 1-x

OpenStudy (amistre64):

if x is rational, then r(r) = x, and none of the others are defined if x is irrational, then i(i) = x, and none of the others are defined